OpenStudy (ms_sp):

In the sophomore class at Dixon High School, the number of students taking French is two-thirds of the number taking Spanish. How many students are studying each language if the total number of students in French and Spanish is 310?

OpenStudy (lannyxx):

so the total number of students is 310 students the total number of students studying french is f, let's say. cuz we don't know yet and the total number of students studying spanish is, yeah s. we know that f + s = 310 students, because that's what the last line tells us. we also know that f = 2/3 * s because it said "The students studying french are 2 thirds the students studying spanish" You can now do a subsitution in f + s = 310, and solve for either f or s then you get both!!!

OpenStudy (ms_sp):

Can you explain it, because it says in the book is "Spanish: 186 and French: 124."

OpenStudy (lannyxx):

well first try to fit that in: s = 186 f = 124 is s + f = 310? is f = 2/3 * s can you confirm that through calculation?

OpenStudy (ms_sp):

124+186 = 310

OpenStudy (lannyxx):

yeah! do you know how to substitute for variables in equations? do you get whay f = 2/3 * s? but because what the equal sign tell you is that what is on the right is exactly what is on the left! f + s = 310. this has two variables, we can't solve it!!!....but what if it had only one? see that f there, replace it with 2/3 * s, no problem since they are the same thing, right? 2/3 * s + s = 310 now you solve for s then once you get s you solve for f f = 310 - s

OpenStudy (ms_sp):

Thank you!

OpenStudy (lannyxx):

yuppie!